### v1

@@ Appendix B Additional Experiments
Figure 11: Using refusal cones to induce refusal across various Qwen 2.5 models with different dimensions. We observe that inducing refusal is generally easier than executing an attack. In this setting, nearly all dimensions maintain strong performance in eliciting refusal responses, even for benign requests.
Figure 12: Attack success rates in refusal cones of different dimensions for Gemma 2 2B where the basis vectors are trained to be representationally independent.
@@
Moving on to the ablation study, Figure 15 presents an analysis of the relationship between the retain loss weight and model performance on the Qwen 2.5 3B model. The left plot illustrates the performance under directional ablation and activation subtraction, with results averaged over five runs per loss weight. For this model, a loss weight of 1 or lower yields the best generalization, while increasing the penalty for unintended side effects on harmless instructions significantly degrades performance. On the right, we examine how the retain loss weights generalize to the validation KL score. This allows us to abstract from specific training conditions and evaluate how effectively the loss weights transfer beyond the training setup.
Figure 15: The left plot shows the relationship between the retain loss weight and performance when using the trained direction for directional ablation and activation subtraction on the Qwen 2.5 3B model, with mean and standard deviation over 5 runs per loss weight. For this specific model, a loss weight of 1 or smaller results in the best generalization, and performance degrades significantly as the direction is penalized more for unintended side-effects on harmless instructions. On the right, we also show performance depending on how the directions generalized to the validation KL-score.
Finally, we assess the performance of the baseline across different (layer, token) combinations. Figure 16 visualizes the effectiveness of the direction selection algorithm from Arditi et al. (2024) for DIM directions in the Qwen 2.5 7B model. Among the evaluated token and layer pairs, only one direction is found to be effective for both inducing refusal through activation addition and maintaining low side effects. Transparent data points indicate (layer, token) combinations that were filtered out due to their inability to induce refusal reliably. Additionally, the red line represents the KL-divergence threshold, used to estimate potential side effects of directional ablation on harmless instructions.
Appendix C Assets

### v2

@@ B.1 Activation Addition
B.2 Directional Ablation on Additional Datasets
For a more robust evaluation of how the RDO directions compare to DIM in terms of performance, we additionally evaluate directional ablation ASR on StrongREJECT(Souly et al., 2024) and the SORRY-Bench base dataset (Xie et al., 2025). We include baseline performance here without any intervention. The results are shown in Figure 11, Figure 12, and Figure 13, and confirm that our findings transfer across datasets.
Figure 11: Attack success rates of refusal directions on StrongREJECT.
Figure 12: Attack success rates of refusal directions on SORRY-Bench.
Figure 13: Attack success rates of refusal directions on JailbreakBench.
B.3 Benchmarks
In Table 1 we showed that RDO directions have significantly lower side-effects on model performance as measured via the reduction in TruthfulQA score. Here, we show the change in benchmark scores for more benchmarks, specifically ARC (Clark et al., 2018), GSM8K (Cobbe et al., 2021), and MMLU (Hendrycks et al., 2020). In Table 4, we see that for most combinations of model and benchmark, the RDO direction has less impact on benchmark score, which confirms the effectiveness of our retain loss and that the RDO directions manipulate refusal more precisely. Finally, we summarize all results in Table 5.
Table 4: Sides effect measured by comparing model performance on benchmark datasets when ablating with either DIM or RDO. Ablating RDO has significantly lower side effects for most models and benchmarks.
Dataset Model DIM RDO Change
TruthfulQA_MC2 Gemma 2 2B 8.0% 4.4% (-3.6%)
Gemma 2 9B 8.3% 4.4% (-3.9%)
Llama 3 8B 4.1% 1.8% (-2.3%)
Qwen 2.5 1.5B 3.6% 2.6% (-1.1%)
Qwen 2.5 7B 4.4% 3.1% (-1.2%)
Qwen 2.5 14B 7.5% 2.5% (-5.0%)
ARC_CHALLENGE Gemma 2 2B 0.2% 0.2% (-0.0%)
Gemma 2 9B 0.6% 0.3% (-0.3%)
Llama 3 8B 1.0% 0.2% (-0.9%)
Qwen 2.5 1.5B 0.7% 0.3% (-0.4%)
Qwen 2.5 7B 0.9% 0.0% (-0.9%)
Qwen 2.5 14B 1.5% 0.6% (-0.9%)
GSM8K Gemma 2 2B 0.6% 0.4% (-0.2%)
Gemma 2 9B 0.1% 0.3% (+0.2%)
Llama 3 8B 1.0% 2.7% (+1.7%)
Qwen 2.5 1.5B 1.1% 0.0% (-1.1%)
Qwen 2.5 7B 0.8% 0.7% (-0.2%)
Qwen 2.5 14B 1.5% 0.5% (-1.1%)
MMLU Gemma 2 2B 0.3% 1.2% (+1.0%)
Gemma 2 9B 1.3% 0.1% (-1.2%)
Llama 3 8B 2.6% 0.1% (-2.5%)
Qwen 2.5 1.5B 1.5% 0.5% (-1.0%)
Qwen 2.5 7B 0.7% 0.1% (-0.6%)
Qwen 2.5 14B 1.2% 0.3% (-0.9%)
Table 5: Comparing RDO and DIM in terms of jailbreaking effectiveness and how jailbreaking affects general capability benchmarks. Each cell in the jailbreaking section contain pairs of ASRs: the first for directional ablation and the second for activation subtraction. The values in the general capability section are computed under directional ablation of the respective directions.
Jailbreaking General Capability
JailbreakBench StrongREJECT SORRY-Bench MMLU ARC-C GSM8K TruthfulQA
ASR $\uparrow$ ASR $\uparrow$ ASR $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$ Acc $\uparrow$
Gemma 2 2B 2.5 1.0 10.4 30.5 42.7 52.3 55.8
  DIM 78.6 / 71.2 80.7 / 74.6 71.0 / 66.3 30.3 42.5 52.9 47.8
  RDO 79.9 / 69.5 80.5 / 71.4 69.5 / 63.3 29.3 42.8 52.7 51.4
Gemma 2 9B 0.3 0.6 7.8 34.8 52.6 74.7 61.1
  DIM 79.1 / 68.3 80.2 / 82.7 69.1 / 69.0 36.1 53.2 74.6 52.8
  RDO 76.8 / 73.0 77.7 / 75.4 68.1 / 66.3 34.7 52.2 75.0 56.7
Llama 3 8B 2.9 1.2 14.0 58.1 48.6 63.5 52.8
  DIM 79.7 / 74.1 81.5 / 81.4 73.0 / 73.4 55.5 47.6 62.5 48.7
  RDO 80.3 / 79.0 83.8 / 84.7 74.4 / 73.2 58.2 48.5 66.2 51.0
Qwen 2.5 1.5B 0.4 3.1 9.2 58.3 38.5 56.4 46.5
  DIM 53.1 / 55.0 65.2 / 60.4 52.5 / 53.0 56.8 37.8 55.3 42.9
  RDO 53.5 / 65.2 61.3 / 71.3 49.1 / 60.1 57.8 38.7 56.4 44.0
Qwen 2.5 3B 8.4 6.9 16.7 64.6 42.8 61.2 57.2
  DIM 68.5 / 54.9 72.2 / 64.1 64.8 / 55.6 64.5 42.7 60.3 54.2
  RDO 67.6 / 54.8 73.6 / 55.9 63.6 / 55.0 64.7 41.6 59.5 54.5
Qwen 2.5 7B 9.1 7.1 22.7 68.8 45.6 77.9 63.1
  DIM 69.2 / 71.0 68.8 / 74.1 63.1 / 64.8 68.1 46.5 77.0 58.7
  RDO 69.3 / 72.0 70.0 / 74.8 63.4 / 66.7 68.7 45.6 77.2 60.0
Qwen 2.5 14B 4.9 2.9 17.9 76.9 52.8 81.7 70.8
  DIM 76.2 / 66.6 80.8 / 70.7 69.6 / 60.0 75.7 51.4 80.2 63.4
  RDO 75.7 / 75.6 79.9 / 76.2 69.0 / 66.2 76.6 52.2 81.3 67.9
Appendix C Ablation Studies
We conduct ablation studies to determine the importance of the three losses in our RDO algorithm.
C.1 Addition and Ablation Loss
We first study how the addition and ablation loss should be balanced. We experiment with the Llama-3-8B-Instruct model and range the loss weights $\lambda_{\text{abl}}$ and $\lambda_{\text{add}}$ from 0 to 1, setting $\lambda_{\text{abl}}=1-\lambda_{\text{add}}$ to balance the weights. We then evaluate attack success rates for both directional ablation and activation subtraction interventions. Figure 14 shows that both loss components are essential for optimal performance. ASR is similar across the 0.2–0.8 weight range, where both methods maintain consistently high attack success rates above 80%, and choosing only one of the losses reduces performance significantly. This finding indicates that while including both loss terms is critical, the precise weight allocation within this range has minimal impact on effectiveness. This robustness simplifies hyperparameter tuning in practice, as practitioners can select any weight configuration within this range without substantially affecting performance.
Figure 14: Ablation study of loss weights $\lambda_{\text{abl}}$ and $\lambda_{\text{add}}$ for Llama-3-8B-Instruct. We compare attack success rates for directional ablation and activation subtraction across different weight balances ($\lambda_{\text{abl}}=1-\lambda_{\text{add}}$). Both methods perform well in the 0.2–0.8 range, with severe degradation at extremes, particularly for directional ablation using only the addition loss (20% ASR at $\lambda_{\text{add}}=1$).
C.2 Retain Loss
Regarding the impact of using the retain loss to minimize side-effects when intervening with our refusal directions, we conduct an ablation study of the retain loss weight $\lambda_{\text{ret}}$ for Qwen2.5-3B-Instruct. We fix the ablation and addition loss weights to their default values (see Table 3) and systematically vary $\lambda_{\text{ret}}$. Figure 15 presents a Pareto analysis plotting ASR under directional ablation against the average reduction in benchmark performance that results from directional ablation. In more detail, the x-axis represents the average change in benchmark scores when intervening with the refusal direction across GSM8K, MMLU, ARC-Challenge, and TruthfulQA relative to the model’s baseline performance (no intervention), while the y-axis shows the corresponding JailbreakBench ASR under directional ablation. The ideal refusal direction would maximize ASR while maintaining or improving benchmark performance (e.g., by preventing the model from inappropriately refusing legitimate questions).
For this specific model, retain weights up to $\lambda_{\text{ret}}=4$ increasingly reduce side-effects on benchmark performance with only marginal ASR reduction. However, beyond this threshold, the ASR drops drastically, indicating that excessive weighting of the retain loss impedes the learning of effective refusal directions.
Figure 15: Ablation study of retain loss weight $\lambda_{\text{retain}}$ for Qwen2.5-3B-Instruct. We plot JailbreakBench ASR versus benchmark performance change (x-axis, averaged over GSM8K, MMLU, ARC-Challenge, TruthfulQA) when intervening with the direction compared to baseline. Higher retain weights improve benchmark preservation with minimal ASR loss up to $\lambda_{\text{retain}}=4$, after which ASR drops significantly. Multiple hyperparameter choices Pareto-dominate the DIM baseline, demonstrating robust improvements over current state-of-the-art.
Importantly, we observe that multiple hyperparameter configurations achieve Pareto dominance over the DIM baseline across both metrics. This demonstrates that our method provides robust improvements over the current state-of-the-art, rather than gains limited to specific hyperparameter choices.
Appendix D Flexibility of Algorithms
We find that DIM can struggle to identify refusal directions with sufficiently low side-effects (according to the heuristic used by Arditi et al. (2024). Figure 16 visualizes the effectiveness of the direction selection algorithm from Arditi et al. (2024) for DIM directions in the Qwen 2.5 7B model. Among the evaluated token and layer pairs, only one direction is found to be effective for both inducing refusal through activation addition and maintaining low side effects. Transparent data points indicate (layer, token) combinations that were filtered out due to their inability to induce refusal reliably. Additionally, the red line represents the KL-divergence threshold, used to estimate potential side effects of directional ablation on harmless instructions.
Appendix E Extended Results for Refusal Cones
In Figure 17 we show the refusal cones for the Gemma model family.
@@ Appendix E Extended Results for Refusal Cones
In Figure 18 we measure the performance of the refusal cones in the Qwen 2.5 model family for activation addition. The models tend to support higher dimensional cones compared to Figure 4, revealing that inducing refusal is significantly easier than disabling refusal via directional ablation. Notably, most directions even in high-dimensional cones remain effective at inducing refusal responses.
Figure 18: Using refusal cones to induce refusal across various Qwen 2.5 models with different dimensions. We observe that inducing refusal is generally easier than executing an attack. In this setting, nearly all dimensions maintain strong performance in eliciting refusal responses.
@@
Appendix F Orthogonal Refusal Directions
In Figure 6 we showed how the orthogonal $\textnormal{RDO}_{\perp}$ interacts with the DIM direction. Here, we show the result for two additional models, on a larger dataset (SORRY-Bench). Figure 22 supports that the DIM direction is greatly influenced by ablating orthogonal refusal directions, supporting our claims in Section 6.1.
Figure 21: Interaction of orthogonal refusal directions directions under directional ablation, measured in terms of cosine similarity to model activations at the last token position on harmful instructions from SORRY-Bench, across different models.
Appendix G Over-refusal
We evaluate the trade-off between over-refusal and safety by analyzing activation addition of the RDO and DIM directions across different strength configurations for the Gemma 2 2B model. Figure 22 shows results for activation addition strengths ($\alpha$) ranging from 0 to $v||_{2}$, where we compute refusal scores on two distinct datasets: harmful instructions from SORRY-Bench and harmless instructions from XSTest (Röttger et al., 2023). Higher scores on SORRY-Bench indicate safer responses to genuinely harmful prompts, while lower scores on XSTest indicate reduced over-refusal on benign inputs that are designed to measure inappropriate refusal behavior. Our method consistently provides better or equivalent trade-offs compared to DIM across all strength configurations, suggesting that RDO directions may be more suited for increasing safety at the same rate of over-refusal compared to DIM.
Figure 22: Trade-off between over-refusal and safety for different activation addition strengths of our refusal directions.
Appendix H Assets
@@ H.2 Datasets
StrongREJECT Souly et al. (2024) Link MIT License
SORRY-Bench Xie et al. (2025) Link Custom License
XSTest Röttger et al. (2023) Link $C$-BY-4.0
@@
MMLU Hendrycks et al. (2020) Link MIT License
ARC Clark et al. (2018) Link $C$-BY-SA-4.0
GSM8K Cobbe et al. (2021) Link MIT License